# FDSG: Forecasting Dynamic Scene Graphs

### v1

@@ Abstract
Dynamic scene graph generation extends scene graph generation from images to videos by modeling entity relationships and their temporal evolution. However, existing methods either generate scene graphs from observed frames without explicitly modeling temporal dynamics, or predict only relationships while assuming static entity labels and locations. These limitations hinder effective extrapolation of both entity and relationship dynamics, restricting video scene understanding. We propose Forecasting Dynamic Scene Graphs (FDSG), a novel framework that predicts future entity labels, bounding boxes, and relationships, for unobserved frames, while also generating scene graphs for observed frames. Our scene graph forecast module leverages query decomposition and neural stochastic differential equations to model entity and relationship dynamics. A temporal aggregation module further refines predictions by integrating forecasted and observed information via cross-attention. To benchmark FDSG, we introduce Scene Graph Forecasting, a new task for full future scene graph prediction. Experiments on Action Genome show that FDSG outperforms state-of-the-art methods on dynamic scene graph generation, scene graph anticipation, and scene graph forecasting. Codes will be released upon publication.
@@ 1 Introduction
Figure 1: Comparison between existing tasks and our FDSG framework. (a) DSGG generates scene graphs and entity bounding boxes from observed video. (b) SGA predicts scene graphs without bounding boxes for future unobserved video. In contrast, FDSG forecasts both scene graph (b) and predicts entity labels and bounding boxes (c) in the future, while also aggregating observed-frame information to enhance both generation and forecasting (d). Our FDSG employs a one-stage end-to-end design, without explicit object tracking, and incorporates forecasting error as an auxiliary loss, enabling joint learning of DSGG and SGA.
@@
  • We propose Scene Graph Forecasting (SGF), a new task that forecasts complete future scene graphs – including entities, bounding boxes, and interactions – from a video stream. This new task complements the existing DSGG and SGA tasks, providing a comprehensive evaluation for extrapolation-based scene graph generation.
@@
  • We rigorously validate our proposed model with the Action Genome dataset [16] on DSGG, SGA, and SGF tasks. Our method achieves superior performance on the SGF task and outperforms the existing methods on both DSGG and SGA tasks, demonstrating that learning to extrapolate can benefit the model’s overall ability of dynamic scene understanding.
@@ 2 Related Work
Figure 2: Our FDSG Framework carries out Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) simultaneously to fully exploit temporal dependencies. Video features are extracted using a CNN backbone and a DINO encoder. Then, (a) Scene Graph Generation Module generates scene graph representations for each observed frame. Subsequently, (b) Forecast Module forecasts a scene graph from a reference frame to the target frame. Finally, (c) Temporal Aggregation Module aggregates information from DSGG and SGF to jointly enhance model performance.
@@
Dynamic Scene Graph Generation. The Dynamic Scene Graph Generation (DSGG) task aims to generate a scene graph for each video frame, focusing on modeling the temporal dynamics of entities and their pairwise relationships. Early static [43, 45, 26, 39, 27] and dynamic scene graph generation [9, 17, 40, 42] models rely on off-the-shelf object detectors such as Faster-RCNN [34], which are also known as two-stage methods. Once entity representations from each frame are obtained, these two-stage methods typically utilize dedicated self-attention [9] or cross-attention layers [17, 40, 42] for temporal feature aggregation. With the recent success of the Transformer-based object detection model DETR [2], more attention has been paid to DETR-based one-stage scene graph generation for images [23, 10, 15] and videos [48, 41]. These methods use a fixed set of queries for entity detection, facilitating finding the correspondence between the target and reference entities or triplets. These methods, although having achieved remarkable success in benchmark datasets, they utilize both the target frame and the reference frames to interpolate scene graphs from observed data. As a result, these models lack the capacity to predict future video frames accurately enough to support dynamic scene graph forecasting.
@@
Anticipating Visual Relationships. Modeling relationships dynamics across frames in scene graphs is crucial for DSGG. The first exploration of anticipating future visual relationships comes from APT [24], which introduces anticipatory pertaining to predict predicates (relationships) in the next frame using reference frames, then fine-tuned on all frames. While this improves the performance on DSGG, APT relies on conventional metrics and does not explicitly evaluate anticipation. More recently, Peddi et al. [33] formalized a Scene Graph Anticipation (SGA) task with a prediction model called SceneSayer, which combines two-stage DSGG models STTran [9] and DSG-DETR [11] and adopts a Neural Differential Equation-based dynamics model to predict future relationships. It focuses on the correctness of predicted visual relationships in the scene graph, assuming that the entity labels and locations remain unchanged in the future. The new evaluation metrics in SGA focus solely on predicting correct relationships, assuming fixed entity labels and locations. However, this overlooks entity dynamics, limiting anticipatory capability to predicting relationships under rigid location and label constraints. Unlike the above methods with partial forecasting, our proposed Scene Graph Forecasting (SGF) task extends the relationship prediction to include both entity locations and labels. This allows for a more comprehensive modeling of motion and relationship dynamics across frames. Moreover, we adopt a one-stage end-to-end method to achieve DSGG and SGF simultaneously with mutual benefits.
@@
Anticipating Entity Labels and Locations. Anticipating entity labels and locations have been explored in various domains. For example, Object Forecasting [37, 19] aims to predict future bounding boxes of tracked objects, which is closely related to object tracking [1]. [37, 19] use RNNs to model entity evolution which rely heavily on tracked object trajectory and low-level features from optical flow estimation [44], but ignore information from high-level scene understanding, such as object interactions. Action Anticipation, on the other hand, aims to predict what is going to happen in the scene given a portion of the activity video [12, 35, 13, 31]. It requires video understanding from an overview perspective. Recently, a novel task, known as Future Action Localization [8], aims to predict the bounding boxes for the central action-taking person(s) in videos from a third-person view. The proposed model AdamsFormer [8] models the dynamics of the central person using a Neural Ordinary Differential Equation model [5] with features extracted from a CNN-based video encoder. However, this task overlooks modeling other objects interacting with the central person. In comparison, our proposed method for forecasting dynamic scene graphs takes a more comprehensive perspective, which includes anticipating relationships of interactions between various entities, entity labels, and locations. This forecasting process requires modeling both low-level features for entity prediction and localization, and high-level understanding of interaction dynamics between entities within the scene.
@@ 3.1 Problem Formulation
Scene Graph Forecasting (SGF): with the same observation time horizon as SGA, but the model predicts both the evolution of predicates and future entity labels and bounding boxes in the anticipatory time horizon. The task of SGF is illustrated in Fig.3.
@@ 3.2 Overview
The overall workflow of our proposed method is illustrated in Fig. 2. Our Scene Graph Generation Module (Fig. 2(a)) generates static scene graphs for each input frame. The Forecast Module (Fig. 2(b)) forecasts scene graphs in the future. The Temporal Aggregation Module (Fig. 2(c)) aggregates temporal information from observed video and from scene graph forecasts, and updates the final prediction for the target frame.
@@ 3.3 Scene Graph Generation Module
To form triplet representations $z(i,j)$, we follow common practice in the literature [9, 33, 15] and concatenate the output content representation $\hat{o}^{c}(i)$ of the subject with the content representation $\hat{o}^{c}(j)$ of all other objects. The predicate is implicitly encoded:
@@
Finally, the scene graph classification head includes a linear layer to classify entity labels, a three-layer Multi-Layer Perceptron (MLP) for bounding box regression, and a two-layer MLP for predicate classification.
@@ 3.4 Forecast Module
Our Forecast Module consists of a Triplet Dynamics Model and a Location Dynamics Model (LDM). In the Triplet Dynamics Model, we utilize a NeuralSDE (see Supplementary B.3 for more details) to forecast triplet representations from state $z_{T_{0}}$, where $T_{0}\in\{1,...,T\}$, to state $z_{T_{0}+\Delta T}$ at $\Delta T$ time steps later, inspired by SceneSayer [33]:
$z_{T_{0}+\Delta T}=\text{NeuralSDE}(z_{T_{0}},\Delta T).$ (2)
To obtain forecasts for entity contents, we reverse the concatenation operation in Eq. (1) and split the forecasted triplet representation:
@@
Our LDM predicts entity location representations:
@@
While intuitively, another NeuralSDE can be adopted for the LDM, in our experiments, we found that the LDM can be reduced to an elegantly effective Identity Mapping. More detailed ablation study on this can be found in Table 5.
The classification head for forecasted scene graphs takes the same form as in the Scene Graph Generation module:
@@
In the case where the LDM is an Identity Mapping, $\hat{o}^{b}_{T_{0}+\Delta T}=\hat{o}^{b}_{T_{0}}$, and Eq. (6) reduces to
$\hat{b}_{T_{0}+\Delta T}=\sigma(\,\text{MLP}(\hat{o}_{T_{0}+\Delta T}^{c}+\sigma^{-1}(\hat{o}_{T_{0}}^{b})\,).$ (8)
This means that the anticipated object bounding box $\hat{o}_{T_{0}+\Delta T}^{b}$ is the sum of a reference bounding box from the last observation $\hat{o}_{T_{0}}^{b}$, and a residual term provided by the anticipated object content representation $\text{MLP}(\hat{o}_{T_{0}+\Delta T}^{c}(j))$. The Triplet Dynamics Model in Eq. (2) is therefore responsible for learning such a residual. More details and analysis on the realization of the LDM are presented in Sec. 4.3.
@@ 3.5 Temporal Aggregation Module
For each frame $V_{t}$, SGG Module generates a set of triplet representations $Z_{t}=\{z_{t}(i,j)\,|\,i\in\text{sbj},j\in\text{obj}\}$. To aggregate temporal information for the target frame $V_{T_{g}}$, following [49, 41], we sample reference frames $\{V_{\text{ref}_{1}},...,V_{\text{ref}_{n}}\}$ from the video. Then we use cascaded Transformer decoders to aggregate information, first from the reference triplet representations $Z_{\text{ref}}=\{Z_{\text{ref}_{1}},...,Z_{\text{ref}_{n}}\}$, and then from forecasted representations $Z_{T_{0}+\Delta T}$.
For reference frame aggregation, the triplet representations of the target frame $Z_{T_{g}}$ is updated by first doing self-attention and then cross-attention with reference triplets:
$Z_{T_{g}}=\text{CrossAttn}(\,\text{SelfAttn}(Z_{T_{g}}),\text{TopK}(Z_{\text{ref}},k)\,)$ (9)
where TopK selects $k$ most confident reference triplets in terms of object and predicate classification scores. We use multiple decoders and the value of $k$ is gradually reduced in order to obtain progressively refined features. See Supplementary B.1 for details on attention layers.
To aggregate information from forecasts, we use
@@
where $T_{0}$ is selected from a reference frame $\text{ref}_{n}$ and $T_{0}+\Delta T=T_{g}$. As shown in [8, 33], a larger anticipatory window benefits NeuralSDE performance; hence, we select $\text{ref}_{n}$ such that the time difference $\Delta T$ is as large as possible. A more detailed analysis can be found in [33].
We first aggregate information from reference frames with lower uncertainty, followed by forecasts with higher uncertainty. Alternative aggregation strategies underperform, and details are provided in Supplemantary C.4.

### v2

@@ Abstract
Dynamic scene graph generation extends scene graph generation from images to videos by modeling entity relationships and their temporal evolution. However, existing methods either generate scene graphs from observed frames without explicitly modeling temporal dynamics or predict only relationships while assuming static entity labels and locations. These limitations hinder effective extrapolation of both entity and relationship dynamics, restricting video scene understanding. We propose Forecasting Dynamic Scene Graphs (FDSG), a novel framework that predicts future entity labels, bounding boxes, and relationships, for unobserved frames while also generating scene graphs for observed frames. Our scene graph forecast module leverages query decomposition and neural stochastic differential equations to model entity and relationship dynamics. A temporal aggregation module further refines predictions by integrating forecasted and observed information via cross-attention. To benchmark FDSG, we introduce Scene Graph Forecasting, a new task for full future scene graph prediction. Experiments on Action Genome show that FDSG outperforms state-of-the-art methods on dynamic scene graph generation, scene graph anticipation, and scene graph forecasting. Codes will be released upon publication.
Index Terms:
Scene Understanding, Scene Graph Generation, Dynamic Scene Graph, Scene Graph Forecasting
@@ 1 Introduction
Figure 1: Comparison between existing tasks and our FDSG framework. (a) DSGG generates scene graphs and entity bounding boxes from observed video. (b) SGA predicts scene graphs without bounding boxes for future unobserved video. In contrast, FDSG forecasts complete scene graphs, including both relationships (b) and entity labels and bounding boxes (c) in the future. FDSG also aggregates observed-frame information to enhance both generation and forecasting (d). Our FDSG employs a one-stage end-to-end design, without explicit object tracking, and incorporates forecasting error as an auxiliary loss, enabling joint learning of DSGG and SGA.
@@
  • We propose Scene Graph Forecasting (SGF), a new task that forecasts complete future scene graphs – including entities, bounding boxes, and interactions – from a video stream. This new task complements the existing DSGG and SGA tasks, providing a comprehensive evaluation for the extrapolation-based scene graph generation.
@@
  • We rigorously validate our proposed model with the Action Genome dataset [18] on DSGG, SGA, and SGF tasks. Our method achieves superior performance on the SGF task and outperforms the existing methods on both DSGG and SGA tasks, demonstrating that learning to extrapolate can benefit the model’s overall ability for dynamic scene understanding.
@@ 2 Related Work
Figure 2: Our FDSG Framework carries out Dynamic Scene Graph Generation (DSGG) and Scene Graph Forecasting (SGF) simultaneously to fully exploit temporal dependencies. Video features are extracted using a CNN backbone and a DINO encoder. Then, (a) Scene Graph Generation Module generates scene graph representations for each observed frame. Subsequently, (b) Forecast Module forecasts a scene graph from a reference frame (gray representations) to the target frame (orange representations). Finally, (c) Temporal Aggregation Module aggregates information from DSGG and SGF to jointly enhance model performance. The output enhanced scene graph representations are shown in dark green.
2.1 Dynamic Scene Graph Generation
@@
The Dynamic Scene Graph Generation (DSGG) task aims to generate a scene graph for each video frame, focusing on modeling the temporal dynamics of entities and their pairwise relationships. Early static [19, 20, 21, 22, 23] and dynamic scene graph generation [6, 7, 8, 9] models rely on off-the-shelf object detectors such as Faster-RCNN [24], which are also known as two-stage methods. Once entity representations from each frame are obtained, these two-stage methods typically utilize dedicated self-attention [6] or cross-attention layers [7, 8, 9] for temporal feature aggregation. With the recent success of the Transformer-based object detection model DETR [25], more attention has been paid to DETR-based one-stage scene graph generation for images [26, 27, 28] and videos [10, 11]. These methods use a fixed set of queries for entity detection, facilitating the finding of the correspondence between the target and reference entities or triplets. These methods, although having achieved remarkable success in benchmark datasets, they utilize both the target frame and the reference frames to interpolate scene graphs from observed data. As a result, these models lack the capacity to predict future video frames accurately enough to support dynamic scene graph forecasting.
2.2 Anticipating Visual Relationships
@@
Modeling relationship dynamics across frames in scene graphs is crucial for DSGG. The first exploration of anticipating future visual relationships comes from APT [14], which introduces anticipatory pertaining to predicting predicates (relationships) in the next frame using reference frames, then fine-tuned on all frames. While this improves the performance on DSGG, APT relies on conventional metrics and does not explicitly evaluate anticipation. More recently, Peddi et al.[15] formalized a Scene Graph Anticipation (SGA) task with a prediction model called SceneSayer, which combines two-stage DSGG models STTran [6] and DSG-DETR [29] and adopts a Neural Differential Equation-based dynamics model to predict future relationships. It focuses on the correctness of predicted visual relationships in the scene graph, assuming that the entity labels and locations remain unchanged in the future. The new evaluation metrics in SGA focus solely on predicting correct relationships, assuming fixed entity labels and locations. However, this overlooks entity dynamics, limiting anticipatory capability to predicting relationships under rigid location and label constraints. Unlike the above methods with partial forecasting, our proposed Scene Graph Forecasting (SGF) task extends the relationship prediction to include both entity locations and labels. This allows for a more comprehensive modeling of motion and relationship dynamics across frames. Moreover, we adopt a one-stage end-to-end method to achieve DSGG and SGF simultaneously with mutual benefits.
2.3 Anticipating Entity Labels and Locations
@@
Anticipating entity labels and locations has been explored in various domains. For example, Object Forecasting [30, 31] aims to predict future bounding boxes of tracked objects, which is closely related to object tracking [32]. [30, 31] use RNNs to model entity evolution, which rely heavily on tracked object trajectory and low-level features from optical flow estimation [33], but ignore information from high-level scene understanding, such as object interactions. Action Anticipation, on the other hand, aims to predict what is going to happen in the scene given a portion of the activity video [34, 35, 36, 37]. It requires video understanding from an overview perspective. Recently, a novel task, known as Future Action Localization [38], aims to predict the bounding boxes for the central action-taking person(s) in videos from a third-person view. The proposed model, AdamsFormer [38], models the dynamics of the central person using a Neural Ordinary Differential Equation model [39] with features extracted from a CNN-based video encoder. However, this task overlooks modeling other objects interacting with the central person. In comparison, our proposed method for forecasting dynamic scene graphs takes a more comprehensive perspective, which includes anticipating relationships of interactions between various entities, entity labels, and locations. This forecasting process requires modeling both low-level features for entity prediction and localization, and high-level understanding of interaction dynamics between entities within the scene.
@@ 3.1 Problem Formulation
Scene Graph Forecasting (SGF): with the same observation time horizon as SGA, but the model predicts both the evolution of predicates and future entity labels and bounding boxes in the anticipatory time horizon. Concretely, the model uses DSGG results for the last observed frame $\hat{G}_{T}$ as initial value, and forecasts scene graphs $\{\hat{G}_{t}\}_{t=T+1}^{H}$. Notably, SGF predicts new entities for $\{\hat{G}_{t}\}_{t=T+1}^{H}$, instead of re-using entities in $\hat{G}_{T}$. The task of SGF is illustrated in Fig. 3.
@@ 3.2 Overview
The overall workflow of our proposed method is illustrated in Fig. 2. Our Scene Graph Generation Module (Fig. 2(a)) generates static scene graphs for each input frame. The Forecast Module (Fig. 2(b)) forecasts scene graphs in the future. The Temporal Aggregation Module (Fig. 2(c)) fuses temporal information from observed video and scene graph forecasts, and updates the final prediction for the target frame.
@@ 3.3 Scene Graph Generation Module
To form triplet representations $z(i,j)$, we follow common practice [6, 15, 28] and concatenate the output content representation $\hat{o}^{c}(i)$ of the subject with the content representation $\hat{o}^{c}(j)$ of all other objects. The predicate is implicitly encoded within the triplet:
@@
Finally, the scene graph classification head includes a linear layer to classify entity labels using $\hat{o}^{c}$, a three-layer Multi-Layer Perceptron (MLP) for bounding box regression from $\hat{o}^{b}$ and $\hat{o}^{c}$, and a two-layer MLP for predicate classification using $z$:
$\hat{c}(j)=\text{Linear}(\hat{o}^{c}(j)),$ (2)
$\hat{b}(j)=\sigma(\,\text{MLP}(\hat{o}^{c}(j))+\sigma^{-1}(\hat{o}^{b}(j))\,),$ (3)
$\hat{p}(i,j)=\text{MLP}(z(i,j)),$ (4)
@@ 3.4 Forecast Module
Our Forecast Module consists of a Triplet Dynamics Model (TDM) and a Location Dynamics Model (LDM).
The TDM takes a holistic view, modeling scene evolution at the triplet level:
$z_{T_{0}+\Delta T}=\text{TDM}(z_{T_{0}},\Delta T).$ (5)
Here, the triplet representation $z$ is formed by concatenating the content representations of the subject and object entities, $o^{c}(i)$ and $o^{c}(j)$, as defined in Eq. (1). To recover individual entity forecasts from the triplet forecast, we reverse this concatenation:
@@
To forecast entity bounding boxes, we incorporate both the location $\hat{o}^{b}_{T_{0}+\Delta T}$ and content $\hat{o}_{T_{0}+\Delta T}^{c}$ of each forecasted entity (Eq. (9)). As the contribution from the entity content $\hat{o}_{T_{0}+\Delta T}^{c}$ is modeled by the TDM (Eqs. (5), (6)), the Location Dynamics Model (LDM) aims to predict the contribution from the entity location term:
@@
The classification head for forecasted scene graphs takes the same form as in the Scene Graph Generation module Eqs. (2 - 4).
@@
3.4.1 Triplet Dynamics Model (TDM) Realization
We implement the TDM using a NeuralSDE [17], which forecasts both the triplet representation $z$ and the associated entity content representations $\hat{o}^{c}$. A NeuralSDE models an initial value problem with neural network–parameterized drift and diffusion terms:
$\begin{split}&\text{NeuralSDE}(z_{T_{0}},\Delta T;\theta_{c},\phi_{c},W_{c})\\ &=z_{T_{0}}+\int_{T_{0}}^{T_{0}+\Delta T}\mu_{\theta_{c}}(z_{T_{0}})\mathrm{d}t\\ &+\int_{T_{0}}^{T_{0}+\Delta T}\nu_{\phi_{c}}(z_{T_{0}})\mathrm{d}W_{c}(t).\end{split}$ (11)
where $\mu_{\theta_{c}}(\cdot)$, $\nu{\phi_{c}}(\cdot)$, and $W_{c}(\cdot)$ denote drift terms parameterized by $\theta_{c}$, diffusion terms parameterized by $\phi_{c}$, and a Wiener process, respectively [39].
Based on the above analysis, our TDM forecasts simultaneously the predicates (Eq. (10)), entity labels (Eqs. (6),(8)) and residual term of entity bounding boxes (Eq. (9)):
$\begin{split}z_{T_{0}+\Delta T}(i,j)&=[\hat{o}^{c}_{T_{0}+\Delta T}(i);\hat{o}^{c}_{T_{0}+\Delta T}(j)]\\ &=\text{NeuralSDE}(z_{T_{0}}(i,j),\Delta T;\theta_{c},\phi_{c},W_{c})\end{split}$ (12)
It is worth mentioning that, with our compact triplet representation, the TDM is capable of forecasting not only the predicates but also complete scene graphs, which therefore extends beyond the predicate anticipation module (LDPU) of SceneSayer [15].
3.4.2 Location Dynamics Model (LDM) Realization
We explore several realizations for the LDM, including Identity Mapping, MLP, NeuralODE [39] and NeuralSDE, and report a comparative analysis in the ablation study Sec. 5.3.2. Among these, the Identity Mapping proved to be both simple and effective for representing future entity locations $\hat{o}^{b}_{T_{0}+\Delta T}$, and is therefore adopted as our default LDM.
With Identity Mapping, the predicted location remains unchanged: $\hat{o}^{b}_{T_{0}+\Delta T}=\hat{o}^{b}_{T_{0}}$, and Eq. (9) reduces to
$\hat{b}_{T_{0}+\Delta T}=\sigma(\,\text{MLP}(\hat{o}_{T_{0}+\Delta T}^{c})+\sigma^{-1}(\hat{o}_{T_{0}}^{b})\,).$ (13)
This formulation reveals that the predicted bounding box $\hat{b}_{T_{0}+\Delta T}$ consists of a reference box from the last observed frame $\hat{o}_{T_{0}}^{b}$, combined with a residual predicted from the content representation via $\text{MLP}(\hat{o}_{T_{0}+\Delta T}^{c})$. Consequently, the TDM (Eq. (5)) is responsible for learning this residual.
@@ 3.5 Temporal Aggregation Module
TABLE I: Comparison of evaluation settings for tasks DSGG, SGA and SGF. GT means the output uses ground truth from the input and is therefore not evaluated.
Task Protocol Observation / Prediction Window Input Evaluated Output
Input Reference Output Video Entity Label Entity BBox Entity Label Entity BBox Predicate
DSGG PredCLS 1$\rightarrow$ H 1 $\rightarrow$ H 1$\rightarrow$ H GT GT
SGCLS - GT
SGDET - -
SGA GAGS 1$\rightarrow$ T 1$\rightarrow$ T T+1$\rightarrow$ H $\times$ $\times$
PGAGS -
AGS - -
SGF SGDET 1$\rightarrow$ T 1$\rightarrow$ T T+1$\rightarrow$ H - -
For each frame $V_{t}$, the SGG Module generates a set of triplet representations $Z_{t}=\{z_{t}(i,j)\,|\,i\in\text{sbj},j\in\text{obj}\}$. To aggregate temporal information for the target frame $V_{T_{g}}$, following previous DETR-based models [40, 11], we sample reference frames $\{V_{\text{ref}_{1}},...,V_{\text{ref}_{n}}\}$ from the video. Then we use cascaded Transformer decoders to aggregate information, first from the reference triplet representations $Z_{\text{ref}}=\{Z_{\text{ref}_{1}},...,Z_{\text{ref}_{n}}\}$, and then from forecasted representations $Z_{T_{0}+\Delta T}$.
For reference frame aggregation, the triplet representations of the target frame $Z_{T_{g}}$ are updated by first doing self-attention and then cross-attention with reference triplets:
$Z_{T_{0}}^{(l)}$ $=\text{CrossAttn}(\,\text{SelfAttn}(Z_{T_{0}}^{(l-1)}),Z_{\text{ref}}^{(l)}\,),$ (14)
$Z_{\text{ref}}^{(l)}$ $=\text{TopK}(Z_{\text{ref}}^{(l-1)},k_{l}),$ (15)
where TopK selects $k$ most confident reference triplets in terms of object and predicate classification scores. We use multiple decoders, with $l$ denoting the $l$-th decoder layer, and the value of $k$ is gradually reduced to obtain progressively refined features.
To aggregate information from forecasts $Z_{T_{0}+\Delta T}$, we use
@@
where $T_{g}$ corresponds to the target frame, $T_{0}$ is selected from a reference frame $\text{ref}_{n}$. As shown in [38, 15], a larger anticipatory window benefits NeuralSDE performance; hence, we select $T_{0}$ such that the time difference $\Delta T$ is as large as possible, and the time steps should match that of the target frame: $T_{0}+\Delta T=T_{g}$.
We first aggregate information from reference frames with lower uncertainty, followed by forecasts with higher uncertainty. Alternative aggregation strategies underperform, and details are provided in the ablation study Sec. 5.3.3.